OpenStudy (alli6106):

A table is on sale for 26% off. The sale price is $333 . What is the regular price?

OpenStudy (mertsj):

If the table is 26% off, you have to pay 74% so 74% of some number is $333 Write the equation that says that and solve it.

OpenStudy (alli6106):

.............

OpenStudy (anonymous):

make "some number" x

OpenStudy (anonymous):

The regular price is what it'd cost at 100 %. But what you know is that at 26 % off the 100 %, it costs $333. That means 100 % - 26 % = 74 % equals $333..

OpenStudy (anonymous):

Got any clue how to write the equation?

OpenStudy (alli6106):

kinda

OpenStudy (anonymous):

One way to put it becomes 0.74 * x = 333

OpenStudy (mathstudent55):

Just translate Mertsj's second line into a math equation To change a percent into a decimal, move the decimal point 2 places tot eh left. For example, 54% as a decimal is 0.54 In math, usually "of" means "times". Also, "is" means "=".

OpenStudy (mathstudent55):

Also, "some number" is an unknow, so use "x" for it.

OpenStudy (alli6106):

what is the regular price

OpenStudy (anonymous):

The regular price is x. :) Can you solve 0.74 * x = 333?

OpenStudy (alli6106):

450

OpenStudy (mathstudent55):

English: \(\color{red}{74 \%} ~~~~\color{blue}{of} ~~~~~\color{green}{some ~number} ~\color{brown}{is} ~~~$333\) Math: \(\color{red}{0.74} ~~~~\color{blue}{\times} ~~~~~~~~~~~~~~\color{green}{x} ~~~~~~~~~~~~~\color{brown}{=} ~$333\) Or simply: \(0.74x = 333\) Solve for x by dividing both sides by 0.74